OpenStudy (anonymous):

A number added to itself twelve times is five more than twice the number. What is the number?

OpenStudy (nory):

Write it algebraically: x + 12x = 5 + 2x Does it make sense how I got that?

OpenStudy (anonymous):

X=5/11

OpenStudy (nory):

@ranjeet619 , you should tell how you got that.